The Marginal Cost for producing x Tshirts is modeled by the function C''(x)=3x-7 . Find the cost function if the fixed cost is 350$.
Answer #1


C(x) = 3x - 7 The antiderivative of the marginal cost is C(x) dx = |(34– 7)dx+C C(x) = 3 -7x+C C(7)=*-7z+C given that, C(O)